OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (-8, 0), (1, 5)

OpenStudy (aum):

Slope m of the line = (y2-y1) / (x2-x1) = (5-0) / (1- -8) = 5/9 y = mx + b y = 5/9x + b It passes through the point (1,5) 5 = 5/9 * 1 + b b = 5 - 5/9 = 40/9 y = 5/9x + 40/9 This is the slope-intercept form. Convert it to standard form: multiply by 9: 9y = 5x + 40 subtract 9y: 0 = 5x - 9y + 40 subtract 40: -40 = 5x - 9y or 5x - 9y = -40

OpenStudy (anonymous):

thanks